A boy is born when his father is 27 years old. How old will he be when he is exactly half his fathers' age?
Found 2 solutions by Edwin McCravy, TamiSrcKa:
Answer by Edwin McCravy(20056)
Answer by TamiSrcKa(4)

father 27
son 0
...
father 2x
son x
_______
Solution:
27-0=2x-x
27=x
>>>The son will be 27, when his father will be 2x=2*27=54.